Solve :


P² -2xyp + 4y²=0, where p=dy/dx

1
Expert's answer
2022-01-27T16:18:57-0500

Using quadratic formula;

"\displaystyle
p=\frac{-(-2xy)\pm\sqrt{(-2xy)^2-4(1)(4y^2)}}{2(1)}\\
\quad=\frac{2xy\pm\sqrt{4x^2y^2-16y^2}}{2}\\
\quad=\frac{2xy\pm\sqrt{4y^2(x^2-4)}}{2}\\
\quad=\frac{2xy\pm2y\sqrt{x^2-4}}{2}\\
\quad=xy\pm y\sqrt{x^2-4}=y(x\pm\sqrt{x^2-4})\\"

But, "\displaystyle
p=\frac{dy}{dx}\\", thus we have;

"\displaystyle
\frac{dy}{dx}=y(x\pm\sqrt{x^2-4})\\
\Rightarrow\frac{1}{y}\frac{dy}{dx}=x\pm\sqrt{x^2-4}" , by method of separation of variable

Integrating both sides wrt x yields;

"\displaystyle
\int\frac{1}{y}\frac{dy}{dx}dx=\int(x\pm\sqrt{x^2-4})dx\\
\int\frac{1}{y}dy=\int x\ dx\pm\int\sqrt{x^2-4}\ dx\qquad\cdots\cdots\cdots\cdots\cdots\cdots\cdots(1)"


But,

"\displaystyle
\int\sqrt{x^2-4}\ dx=4\int\sinh^2p\ dp" , where x=2coshp

"\displaystyle
=2\int(\cosh 2p-1)\ dp" , since "\displaystyle
\cosh 2p=1+2\sinh^2p"

"\displaystyle
=\sinh 2p-2p=2\sinh p \cosh p=2\times\frac{\sqrt{4-x^2}}{2}\times\frac{x}{2}-2 \cosh^{-1}\left(\frac{x}{2}\right)"

"\displaystyle
=\frac{x\sqrt{x^2-4}}{2}-2\cosh^{-1}\left(\frac{x}{2}\right)\qquad\cdots\cdots\cdots\cdots(2)"

Substituting (2) into (1) yields;

"\displaystyle
\log y=\frac{x^2}{2}\pm\left[\frac{x\sqrt{x^2-4}}{2}-2\cosh^{-1}\left(\frac{x}{2}\right)\right]+C, \text{where C is an arbitrary constant}\\
\Rightarrow y=Ae^{\left(\frac{x^2}{2}\pm\left[\frac{x\sqrt{x^2-4}}{2}-2\cosh^{-1}\left(\frac{x}{2}\right)\right]\right)}, \text{where }A=e^C"